OpenStudy (anonymous):

Please help! Let f(x)=x^2+5x, g(x)=6-2x, h(x)=2/x-3; find the following value. 4 g(-2 x)=?

myininaya (myininaya):

First of all we only have to use g since we are asked to simplify 4g(-2x)

myininaya (myininaya):

We are given \[g(x)=6-2x\] so \[g(-2x)=6-2(-2x)\]

myininaya (myininaya):

\[4g(-2x)=4(6-2(-2x))\]

myininaya (myininaya):

You can simplify this right?